# Linear-Size Ancilla Systems for Logical Measurements in QLDPC Codes

### v1

@@ 3.1 Gauged $X$ Ancilla Systems
We follow the notation used to describe the CKBB scheme in Section 2.2, since the initial part of the construction is the same. Recall that we are measuring an $X$-logical operator $\bar{X}_{M}$ with support $V_{0}\subset\mathcal{V}$ and adjacent $Z$-checks $C_{0}\subset\mathcal{C}^{Z}$. As before, we assume there are no other $X$ operators supported on a subset of $V_{0}$. We again consider the induced Tanner graph of $\bar{X}_{M}$: let $J_{V_{0}}:\mathcal{V}\to V_{0}$ and $J_{C_{0}}:\mathcal{C}^{Z}\to C_{0}$ be projection isometries into the corresponding sets, and let $F$ be the submatrix of $H^{Z}$ defined by $F:=J_{C_{0}}^{\top}H^{Z}J_{V_{0}}$. Then $F:C_{0}\to_{Z}V_{0}$ is the adjacency matrix of the induced Tanner graph.
@@
Now we introduce $L$ additional layers $(C_{1},V_{1}),(C_{2},V_{2}),...,(C_{L},V_{L})$ with vertex roles as in Eq. (3) and the connectivity as in the CKBB system with $F:C_{i}\to_{Z}V_{i}$ for even $i$, $F^{\top}:V_{i}\to_{X}C_{i}$ for odd $i$ and $I$ for connections between layers. So far, if we take $L=2d-1$, we have just constructed the CKBB system again.
The CKBB $X$ measurement scheme is a subsystem code with gauge degrees of freedom. These gauge degrees of freedom complicate the analysis in [CKBB22] and are only dealt with in general by using at least $12d-$ 1 layers. In our work we simply promote these gauge degrees of freedom to stabilizers, resulting in a non-subsystem stabilizer code $\mathcal{G}_{X}$.
@@
The $Z$ gauge operators of the CKBB scheme correspond to the nullspace of $F$, that is, $\mathrm{null}(F)=\{c\text{ s.t. }cF=0\}$. This is because the rows of $F^{\top}:V_{i}\to_{X}C_{i}$ for odd $i$ denote $X$ checks within the layer $(C_{i},V_{i})$, and vectors orthogonal to these rows commute with these $X$ checks. Any particular gauge operator $Z(c)$ has equivalent realizations on any $C_{j}$ for any odd $j$, since we can multiply by $Z$ checks in the layers $C_{i}$ with $i$ even. That is, $Z(c\in C_{j})\mathcal{H}_{Z}(c\in C_{j+1})=Z(cF\in V_{j+1})Z(c\in C_{j+2})=Z(c\in C_{j+2})$.
@@
To find a set of additional $Z$ checks that remove these gauge qubits, we find a matrix $G$ whose rows span $\mathrm{null}(F)$, that is, $GF=0$. Minimizing row and column weight of $G$ minimizes the degree of the Tanner graph. Then we introduce a new set of $Z$ checks $U_{L}$ and connect $G:U_{L}\to_{Z}C_{L}$. This completes the construction of $\mathcal{G}_{X}$. A sketch of $\mathcal{G}_{X}$ is shown in Figure 2.
@@
Figure 2: Diagram of the gauged $X$ ancilla system $\mathcal{G}_{X}$ with $L=5$. The original code $\mathcal{G}$’s Tanner graph is a subgraph of this graph supported on $\mathcal{V},\mathcal{C}^{X},$ and $\mathcal{C}^{Z}$. The first layer $(V_{1},C_{1})$ is called the interface, and the remaining layers and the gauge checks $U_{L}$ are called the module.
@@
We defer the proof of this theorem to Appendix A.1.
Remark 2.
Sometimes the gauge checks in $G$ are already products of other $Z$ stabilizers of $\mathcal{G}_{X}$ and hence do not need to be added as explicit checks in $U_{L}$. This happens for the gross code as we will see in Section 4 and for hypergraph product codes.
@@ 3.2 $X$ Measurement Protocol
We now describe the measurement protocol that allows us to extract the measurement result of $\bar{X}_{M}$ while leaving the remaining logical information undamaged. This is achieved through construction of a decoding hypergraph $\mathcal{D}$, which is a collection of detectors. We measure the checks of either $\mathcal{G}$ or $\mathcal{G}_{X}$ for several rounds. Then, a detector is the parity of a subset of these check measurement outcomes, understood as a hyperedge within $\mathcal{D}$. In this section we consider the noise-free case. Our objective is to demonstrate that, although the check measurement outcomes are sometimes non-deterministic, the detector parities are always zero. We assess the fault-tolerance properties of $\mathcal{D}$ in the next section.
$\mathcal{D}$ splits into steps $\mathcal{D}_{0},...,\mathcal{D}_{R}$, corresponding to $R$ measurements of the checks of $\mathcal{G}_{X}$. The first and final steps $\mathcal{D}_{0}$ and $\mathcal{D}_{d}$ handle the code deformation between $\mathcal{G}$ and $\mathcal{G}_{X}$.
As shown in Figure 2, we call the first layer of $\mathcal{G}_{X}$ ($V_{1},C_{1}$) the interface, and the remaining layers are the module. There is no module when $L=1$, so the gauge checks $U_{1}$ are considered part of the interface. The stabilizers $Z(C_{0})$ and $Z(C_{2})$ have support on the interface qubits $C_{1}$. Let $\tilde{C}_{0}$ and $\tilde{C}_{2}$ be modified versions of these checks that are missing support on $C_{1}$. This is useful since $\mathcal{G}$ has no support on the interface, and hence has the $Z$-stabilizers $Z(\tilde{C}_{0})$.
$\mathcal{D}$ is constructed as follows. A small example is shown in Figure 3. In the construction, a check is called unreliable if, even in the noise-free case, the outcome of the check is random since it doesn’t commute with the stabilizer of the current state. A reliable check is deterministic in the noise-free case, but not necessarily $+1$.
@@
    To initialize:
    1. (a)
      Initialize the interface qubits, namely qubits in $C_{1}$, to |$\ket{0}$.
    2. (b)
      If there is a module ($L>1$), initialize a stabilizer state on the module qubits $C_{2},V_{3},\cdots,V_{L}$. It has the stabilizers $Z(\tilde{C}_{2})$, $Z(C_{i})$ for even $i\geq 4$, $Z(U_{L})$, and the $X(V_{i})$ for odd $i\geq 3$. We refer to these checks as the module stabilizer.
    3. (c)
      Measure the module stabilizers. Additionally, either measure the stabilizers of $\mathcal{G}$, or consider the measurement outcomes from a prior round. If there is no module ($L=1$), then we can still compute gauge checks $Z(U_{1})$ by measuring the interface qubits in the $Z$ basis and computing the appropriate parities.
@@
    (Fusion step) Measure the stabilizers of $\mathcal{G}_{X}$. Of these, the $X(V_{1})$ checks are unreliable, but everything else is reliable. Let $\mathcal{D}_{0}$ contain parities between pairs of reliable checks from step 1 and step 2 only. This involves taking parities between step 1’s $Z(\tilde{C}_{0})$ and step 2’s $Z(C_{0})$, as well as step 1’s $Z(\tilde{C}_{2})$ and step 2’s $Z(C_{2})$.
    The module stabilizers except $Z(\tilde{C}_{2})$ are identical in step 1 and step 2 so these parities are zero. From round 1 to 2, checks $Z(\tilde{C}_{0})$ and $Z(\tilde{C}_{2})$ have gained support on the interface $C_{1}$ and become $Z(C_{0})$ and $Z(C_{2})$. But since the interface qubits are in the |$\ket{0}$ state, the measurement outcome is unaffected, so these parities should be zero as well.
    Output the parity of the $X(V_{1})$ checks as the measurement result, since their product $\mathcal{H}_{X}(V_{1})$ is $\bar{X}_{M}$ in the noiseless case. This value should be subject to decoder corrections.
@@
    Measure the stabilizers of $\mathcal{G}_{X}$ $R-1\geq 0$ more times. Let the detectors $\mathcal{D}_{i}$ for $0<i<R$ be the parities between successive measurements. Since these are all the same checks, the parities should all vanish.
@@
    (Split step) Measure the stabilizers of $\mathcal{G}$. Measure the interface qubits $C_{1}$ in the $Z$ basis. Measure the module stabilizers.
    Checks $Z(\tilde{C}_{0})$, $Z(\tilde{C}_{2})$, and the interface qubit $Z$-measurements $C_{1}$ are all unreliable. However, we can still compute what the outcome of the reliable checks $Z(C_{0})$ and $Z(C_{2})$ would have been by considering the appropriate parities of $Z(\tilde{C}_{0})$, $Z(\tilde{C}_{2})$ with the $Z$-measurements of $C_{1}$. The parities of these values with the measurements of $Z(C_{0})$ and $Z(C_{2})$ of the prior step feature in $\mathcal{D}_{R}$. These are shown as the fusion correction in Figure 3.
    As in step 1, if there is no module then gauge checks $Z(U_{1})$ can still be computed by considering the appropriate parities of the $Z$ measurements of $C_{1}$. $\mathcal{D}_{R}$ also features detectors comparing these parities to the $Z(U_{1})$ measurement from the prior round.
The initialization of the module can be done independently and in advance. The module when viewed as a code itself has no logical qubits — it is a single stabilizer state. Therefore, the module is reusable in future rounds of logical measurements, if we keep it error-free. In other words, the module can be viewed as a resource state that facilitates fault-tolerant logical measurements. A similar module system appeared for a different purpose in [XAP+24].
@@ 3.3 Code Distance and Fault Distance
This observation lets us bound the minimum distance in terms of $L$ and the boundary Cheeger constant.
@@
Let $P=X(t\in\mathcal{V})$ be an $X$-type logical operator of $\mathcal{G}_{X}$. Consider
$P^{\prime}=P\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\mathcal{H}_{X}(v_{j})=X(t+v_{1}\in\mathcal{V})\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}X(v_{j}F^{\top}\in C_{j})\prod_{\begin{subarray}{c}i=3\\ i\text{ odd}\end{subarray}}^{L}X(v_{i}+v_{i-2}\in V_{i-1}),$ (10)
for arbitrary $v_{j}$ indicating multiplication by an arbitrary combination of checks from the ancilla system. The definition of Cheeger’s constant says that $|v_{j}F^{\top}|\geq\beta\min(|v_{j}|,|V_{0}|-|v_{j}|)$, and it follows that
$|P^{\prime}|$ $\geq|t+v_{1}|+\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\beta\min(|v_{j}|,|V_{0}|-|v_{j}|)+\sum_{\begin{subarray}{c}i=3\\ i\text{ odd}\end{subarray}}^{L}|v_{i}+v_{i-2}|,$ (11)
$\geq\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\Bigg{[}\frac{1}{\lceil L/2\rceil}\Bigg{(}|t+v_{1}|+\sum_{\begin{subarray}{c}i=3\\ i\text{ odd}\end{subarray}}^{j}|v_{i}+v_{i-2}|\Bigg{)}+\beta\min(|v_{j}|,|V_{0}|-|v_{j}|)\Bigg{]}$ (12)
where the second inequality follows by removing some fraction of the $|v_{j}+v_{j-2}|$ terms and rearranging. By repeated use of the triangle inequality, the parenthesized term is at least $|t+v_{j}|$, so
$|P^{\prime}|$ $\geq\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\Bigg{[}\frac{1}{\lceil L/2\rceil}|t+v_{j}|+\beta\min(|v_{j}|,|V_{0}|-|v_{j}|)\Bigg{]}$ (13)
$\geq\frac{1}{\lceil L/2\rceil}\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\Bigg{[}|t+v_{j}|+\min(|v_{j}|,|V_{0}|-|v_{j}|)\Bigg{]},$ (14)
with the second inequality making use of $\beta\lceil L/2\rceil\geq 1$.
Figure 4: Notation describing the intersection of supports of objects in the proof of Theorem 6. Here $\mathcal{V}$ is the set of all qubits in the original code $\mathcal{G}$, $V_{0}$ is the support of the logical $X$ operator being measured, $t\in\mathcal{V}$ describes the support of another logical $X$ operator $P$, and $v_{j}$ describes a set of checks from $V_{j}$ multiplying $P$. Letters $a,b_{j},c_{j},d_{j},e_{j}$ denote the sizes of the sets in which they are drawn.
Now to complete the proof of $|P^{\prime}|\geq d$ it is sufficient to show $|t+v_{j}|+\min(|v_{j}|,|V_{0}|-|v_{j}|)\geq d$ for all $j$. Using the notation of Figure 4, we need $(a+b_{j}+d_{j})+\min(c_{j}+d_{j},b_{j}+e_{j})\geq d$ or, equivalently, both
$a+b_{j}+c_{j}+2d_{j}\geq d\quad\text{and}\quad a+2b_{j}+d_{j}+e_{j}\geq d.$ (15)
Note that both $X(t\in\mathcal{V})$ and $X(t\in\mathcal{V})X(V_{0})$ are logical operators of the original code $\mathcal{G}$ and thus each has weight at least $d$. Translated to the notation of Figure 4, this means $a+b_{j}+c_{j}\geq d$ and $a+d_{j}+e_{j}\geq d$, which in turn imply (15).
@@
In particular, $\mathcal{S}^{\prime}$ and $\mathcal{S}_{X}$ are two different ways of fixing the gauge of $\mathcal{U}$ to get stabilizer codes.

### v3

@@ 3.1 Gauge-fixed $X$ Ancilla Systems
We follow the notation used to describe the CKBB scheme in Section 2.2, since the initial part of the construction is the same. Recall that we are measuring an $X$-logical operator $\bar{X}_{M}$ with support $V_{0}\subset\mathcal{V}$ and adjacent $Z$-checks $C_{0}\subset\mathcal{C}^{Z}$. As before, we assume there are no other $X$ operators supported on a subset of $V_{0}$, and consider the induced Tanner graph of $\bar{X}_{M}$ with adjacency matrix $F:C_{0}\to_{Z}V_{0}$.
@@
We introduce $L$ additional layers $(C_{1},V_{1}),(C_{2},V_{2}),...,(C_{L},V_{L})$ with vertex roles as in Eq. (3) and the connectivity as in the CKBB system with $F:C_{i}\to_{Z}V_{i}$ for even $i$, $F^{\top}:V_{i}\to_{X}C_{i}$ for odd $i$ and $I$ for connections between layers.
Attaching such an ancilla system introduces new stabilizers and gauge degrees of freedom to the code, which could reduce the code distance. Specifically, multiplying a logical operator of the original code by new stabilizers or gauge operators could reduce its weight below $d$. In [CKBB22], this issue is resolved by taking $L=2d-1$.
@@
In this work, we promote these gauge degrees of freedom to stabilizers, resulting in a non-subsystem stabilizer code $\mathcal{G}_{X}$. The $Z$ gauge operators of the CKBB scheme correspond to the nullspace of $F$, that is, $\mathrm{null}(F)=\{c\text{ s.t. }cF=0\}$. This is because the rows of $F^{\top}:V_{i}\to_{X}C_{i}$ for odd $i$ denote $X$ checks within the layer $(C_{i},V_{i})$, and vectors orthogonal to these rows commute with these $X$ checks. Any particular gauge operator $Z(c)$ has equivalent realizations on any $C_{j}$ for any odd $j$, since we can multiply by $Z$ checks in the layers $C_{i}$ with $i$ even. That is, $Z(c\in C_{j})\mathcal{H}^{Z}(c\in C_{j+1})=Z(cF\in V_{j+1})Z(c\in C_{j+2})=Z(c\in C_{j+2})$.
@@
To find a set of additional $Z$ checks that remove these gauge qubits, we find a matrix $G$ whose rows span $\mathrm{null}(F)$, that is, $GF=0$. Then we introduce a new set of $Z$ checks $U_{L}$ and connect $G:U_{L}\to_{Z}C_{L}$. Minimizing row and column weight of $G$ minimizes the degrees added to the Tanner graph. This completes the construction of $\mathcal{G}_{X}$. A sketch of $\mathcal{G}_{X}$ is shown in Figure 2.
@@
Figure 2: Diagram of the gauge-fixed $X$ ancilla system $\mathcal{G}_{X}$ with $L=5$. The original code $\mathcal{G}$’s Tanner graph is a subgraph of this graph supported on $\mathcal{V},\mathcal{C}^{X},$ and $\mathcal{C}^{Z}$. The first layer $(V_{1},C_{1})$ is called the interface, and the remaining layers and the gauge checks $U_{L}$ are called the module.
@@
We defer the proof of this theorem to Appendix A.1. We note that without additional knowledge on the structure of the code or the measured operator, we do not have upper bounds on the row or column weight of $G$. While this is an undesirable caveat, we show theoretical evidence in Section 3.4 suggesting that for many important families of QLDPC codes, these gauge checks may not break the LDPC property. In particular, the gauge checks may be products of other $Z$ stabilizers of $\mathcal{G}_{X}$ and hence do not need to be explicitly measured. This happens for the gross code as we will see in Section 4 and for hypergraph product codes, as we will show in Theorem 13.
@@ 3.2 $X$ Measurement Protocol
We now describe the measurement protocol that allows us to extract the measurement result of $\bar{X}_{M}$ while leaving the remaining logical information undamaged. This protocol can be studied as a space-time decoding hypergraph $\mathcal{D}$, made of copies of the codes $\mathcal{G}$ and $\mathcal{G}_{X}$ and detectors in between. Our objective is to demonstrate that, although the check measurement outcomes are sometimes non-deterministic, we can still find a set of deterministically zero detector parities. We consider the noise-free case in this section, and assess the fault-tolerance properties of $\mathcal{D}$ in the next section.
$\mathcal{D}$ splits into steps $\mathcal{D}_{0},...,\mathcal{D}_{R}$, corresponding to $R$ measurements of the checks of $\mathcal{G}_{X}$. The first and final steps $\mathcal{D}_{0}$ and $\mathcal{D}_{d}$ handle the code deformation between $\mathcal{G}$ and $\mathcal{G}_{X}$. As shown in Figure 2, we call the first layer ($V_{1},C_{1}$) of $\mathcal{G}_{X}$ the interface, and the remaining layers are the module. There is no module when $L=1$, so the gauge checks $U_{1}$ are considered part of the interface. We call a check unreliable if, even in the noise-free case, the outcome of the check is random since it doesn’t commute with the stabilizer of the current state. A reliable check is deterministic in the noise-free case, but not necessarily $+1$.
In the context of this section, we use the notation $\mathcal{H}^{Z}(C_{0})$ to denote the collection of checks in $C_{0}$ with qubit support given by $\mathcal{H}^{Z}$, instead of the product of these checks. The stabilizers $\mathcal{H}^{Z}(C_{0})$ and $\mathcal{H}^{Z}(C_{2})$ have support on the interface qubits $C_{1}$. $H^{Z}(C_{0})$ are the stabilizers from $\mathcal{G}$ and therefore have no support on the interface. Let $H_{\text{mod}}^{Z}(C_{2})$ denote $Z$-checks at $C_{2}$ on the module, which also have no support on the interface. These notation helps us keep track of which checks are reliable.
$\mathcal{D}$ is constructed with the following measurement schedule. A small example is shown in Figure 3.
@@
    Initialization: Initialize the interface qubits, namely qubits in $C_{1}$, to |$\ket{0}$. If there is a module ($L>1$), initialize a stabilizer state on the module qubits $C_{2},V_{3},\cdots,V_{L}$. It has the stabilizers $H_{\text{mod}}^{Z}(C_{2})$, $\mathcal{H}^{Z}(C_{i})$ for even $i\geq 4$, $\mathcal{H}^{Z}(U_{L})$, and $\mathcal{H}^{X}(V_{i})$ for odd $i\geq 3$. We refer to these checks as the module stabilizer.
    Measure the module stabilizers. Additionally, either measure the stabilizers of $\mathcal{G}$, or take the measurement outcomes from a prior round. If there is no module ($L=1$), then we can still compute gauge checks $\mathcal{H}^{Z}(U_{1})$ by measuring the interface qubits in the $Z$ basis and computing the appropriate parities.
@@
    Fusion: Measure the stabilizers of $\mathcal{G}_{X}$. Of these, the $\mathcal{H}^{X}(V_{1})$ checks are unreliable, but everything else is reliable. Let $\mathcal{D}_{0}$ contain parities between pairs of reliable checks from step 1 and step 2 only. This involves taking parities between step 1’s $H^{Z}(C_{0})$ and step 2’s $\mathcal{H}^{Z}(C_{0})$, as well as step 1’s $H_{\text{mod}}^{Z}(C_{2})$ and step 2’s $\mathcal{H}^{Z}(C_{2})$. The module stabilizers except $H_{\text{mod}}^{Z}(C_{2})$ are identical in step 1 and step 2 so these parities are zero. From round 1 to 2, checks $H^{Z}(C_{0})$ and $H_{\text{mod}}^{Z}(C_{2})$ have gained support on the interface $C_{1}$ and become $\mathcal{H}^{Z}(C_{0})$ and $\mathcal{H}^{Z}(C_{2})$. But since the interface qubits are in the |$\ket{0}$ state, the measurement outcome is unaffected, so these parities should be zero as well.
    Output the parity of the $\mathcal{H}^{X}(V_{1})$ checks as the measurement result, since their product $\mathcal{H}^{X}(V_{1})$ is $\bar{X}_{M}$ in the noiseless case. This value should be subject to decoder corrections.
@@
    Measure the stabilizers of $\mathcal{G}_{X}$ for $R-1\geq 0$ more times. Let the detectors $\mathcal{D}_{i}$ for $0<i<R$ be the parities between successive measurements. The parities should all be zero.
@@
    Split: Measure the stabilizers of $\mathcal{G}$, the interface qubits $C_{1}$ in the $Z$ basis, and the module stabilizers.
    Checks $H^{Z}(C_{0})$, $H_{\text{mod}}^{Z}(C_{2})$, and the interface qubit $Z$-measurements $C_{1}$ are all unreliable. However, we can still compute what the outcome of the reliable checks $\mathcal{H}^{Z}(C_{0})$ and $\mathcal{H}^{Z}(C_{2})$ would have been by considering the appropriate parities of $H^{Z}(C_{0})$, $H_{\text{mod}}^{Z}(C_{2})$ with the $Z$-measurements of $C_{1}$. The parities of these values with the measurements of $\mathcal{H}^{Z}(C_{0})$ and $\mathcal{H}^{Z}(C_{2})$ of the prior step feature in $\mathcal{D}_{R}$. These are shown as the fusion correction in Figure 3.
    As in step 1, if there is no module then gauge checks $\mathcal{H}^{Z}(U_{1})$ can still be computed by considering the appropriate parities of the $Z$ measurements of $C_{1}$. $\mathcal{D}_{R}$ also features detectors comparing these parities to the $\mathcal{H}^{Z}(U_{1})$ measurement from the prior round.
The initialization of the module can be done independently and in advance. The module when viewed as a code itself has no logical qubits — it is a single stabilizer state. Therefore, the module is reusable in future rounds of logical measurements, if we keep it error-free. In other words, the module can be viewed as a resource state that facilitates fault-tolerant logical measurements. A similar module system appeared for a different purpose in [XBAP+24].
@@ 3.3 Code Distance and Fault Distance
Lemma 4 (Expansion Lemma).
If the induced Tanner graph of $\bar{X}_{M}$ is $F:C_{0}\to V_{0}$, suppose $F^{\top}:V_{1}\to C_{1}$ has boundary Cheeger constant $\beta$. Given a set $s\in V_{0}$, suppose we multiply $X(s)$ by a set of $X$-stabilizers in the ancilla patch, namely $v_{j}\in V_{j}$ for odd $j\in[L]$. Suppose $m=\operatorname*{arg\,min}_{i}\min(|v_{i}|,|V_{0}|-|v_{i}|)$. If $L/2\geq 1/\beta$, then
$|X(s)\cdot\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\mathcal{H}_{X}(v_{j}\in V_{j})|$ $\geq|s+v_{m}|+|v_{m}+v_{L}|+\min(|v_{m}|,|V_{0}|-|v_{m}|),$ (10)
$\geq\min(|s|,|V_{0}|-|s|).$ (11)
Proof.
Here we have
$|X(s)\cdot\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\mathcal{H}_{X}(v_{j}\in V_{j})|$ $=|s+v_{1}|+\sum_{\begin{subarray}{c}i=3\\ i\text{ odd}\end{subarray}}^{L}|v_{i}+v_{i-2}|+\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}|v_{j}F^{\top}|,$ (12)
$\geq|s+v_{1}|+\sum_{\begin{subarray}{c}i=3\\ i\text{ odd}\end{subarray}}^{L}|v_{i}+v_{i-2}|+\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\beta\min(|v_{j}|,|V_{0}|-|v_{j}|),$ (13)
$\geq|s+v_{1}|+\sum_{\begin{subarray}{c}i=3\\ i\text{ odd}\end{subarray}}^{L}|v_{i}+v_{i-2}|+\min(|v_{m}|,|V_{0}|-|v_{m}|),$ (14)
$\geq|s+v_{m}|+\sum_{\begin{subarray}{c}i=m+2\\ i\text{ odd}\end{subarray}}^{L}|v_{i}+v_{i-2}|+\min(|v_{m}|,|V_{0}|-|v_{m}|),$ (15)
$\geq|s+v_{m}|+|v_{m}+v_{L}|+\min(|v_{m}|,|V_{0}|-|v_{m}|).$ (16)
Note that the last two inequalities follow from triangle inequality. Our claim holds by another application of triangle inequality. ∎
@@
Combining the two preceding lemmas, we obtain the following bound on the code distance of the merged code.
@@
Let $P=X(t\in\mathcal{V})$ be an $X$-type logical operator of $\mathcal{G}$. Consider
$P^{\prime}=P\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}\mathcal{H}_{X}(v_{j}).$ (17)
Let $s=t\cap|V_{0}|$, from Lemma 4, we see that $|P^{\prime}|\geq|t|-|s|+\min(|s|,|V_{0}|-|s|)$. If $|s|\leq|V_{0}|/2$, we see that $|P^{\prime}|\geq|t|\geq d$. Otherwise, we have
$|P^{\prime}|\geq|t|-|s|+|V_{0}|-|s|=|t\setminus V_{0}|+|V_{0}\setminus t|\geq d,$
by Lemma 5. We conclude that the $X$ distance of $\mathcal{G}_{X}$ is at least $d$. ∎
As a corollary, we see that the space overhead of our ancilla system is $O(d/\beta)$. We remark that boundary expansion is a demanding property, and for general codes we expect $1/\beta$ to be super-constant. However, in the case where every check in the induced Tanner graph acts on exactly two qubits in the support of $\bar{X}_{M}$, the induce Tanner graph can be interpreted as a simple graph by turning every check into an edge. Boundary Cheeger constant on the induced Tanner graph then translates to the usual Cheeger constant on this simple graph. Since edge expansion is a much more common property, for such codes the space overhead may be linear. We note that in [WY24] and [IGND24], instead of using the induced Tanner graph (which is generally a hypergraph if we treat each induced check as a hyperedge) for measurement, the proposed schemes constructed and used a compatible expander graph with sufficient edge expansion (Cheeger constant $\geq 1$) for measurement. This novelty guarantees a low space-overhead ancilla system for any QLDPC code. We also note that in [SJOY24], the requirement on edge expansion is further relaxed to a notion called relative expansion.
We also consider Theorem 6 a worst case upper bound on the number of layers needed and expect practical cases to be much better. The reasoning behind is simple: stabilizers in the ancilla system may only affect $X$-logical operators which intersect $\bar{X}_{M}$. To provide a worse case analysis, we assume no knowledge of the intersection structure of $X$-operators and consider product of all subsets of stabilizers $v\in V$. In reality, only those subsets that touches another $X$-logical operator needs to be expanding. This is exemplified by our case study on the [[144,12,12]] gross code, where we found operators for which a mono-layer ancilla is distance preserving, see Section 4. It is also supported by results from [Cow24], where it was verified numerically (for the CKBB scheme) that for many small-to-medium codes the number of ancilla layers needed is 5 or less.
@@
In particular, $\mathcal{S}^{\prime}$ and $\mathcal{S}_{X}$ are two different ways of fixing the gauge of $\mathcal{U}$ to get stabilizer subspace codes.